If the long-run total cost curve for each firm is given by TC = 400Q – 10Q2 + Q3, where Q is the quantity of output, in the long run, the marginal cost when Q=10 is , fixed costs at Q=10 are and average variable cost is. READ THE QUESTION CAREFULLY BEFORE RESPONDING.


1
Expert's answer
2020-12-21T03:50:03-0500
"MC=400-20Q+3Q^2"


"MC(Q=10)=400-200+3*100=500"


"TC=FC+VC"


"AVC=\frac {VC}{Q}"


"VC=TC-FC=400Q-10Q^2+Q^3-500"


"AVC=400-10Q+Q^2-\frac {500}{Q}"